# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ D.2 Sliding-window probabilistic extraction procedure
We use $100$-token sequences $\bm{z}$}, with a prefix (prompt) length of $50$ tokens and a (target) suffix length of $50$ tokens. We pick $100$ tokens and a 5050/$50$ split because this is accepted in the literature as a reasonable minimum for being confident that extracted suffixes are reflective of memorization [18, 16, 74, 60]. $50$ tokens for the prefix is less for validity reasons. Of course, this prefix length uses an amount of context that is equivalent to the output, and so is compelling in this respect. However, when we are confident that our measurements are capturing valid instances of extraction, using longer prefixes can help discover more underlying memorization. (Indeed, this is part of where discoverable extraction gets its name; Carlini et al. [18] demonstrate the “discoverability phenomenon” with respect to showing how extraction success can increase with longer-context prompts.) We discuss this further (and include experiments with varied prefix lengths) in Section 5 and Appendix E. Other work uses shorter sequences (e.g., $64$ tokens and a 3232/$32$ token split [9]).
@@
For a given book, we start at the beginning of the text file in Books3. We sample a chunk of text that is sufficiently long to contain $100$ tokens of corresponding tokenized text: to do this, we take a chunk of $800$ characters, tokenize, take the first $100$ tokens of the resulting tokenized sequence as the sequence, and discard the rest. (Each $100$ token sequence is typically, but not always, reflective of 300300-$400$ characters of text. Rare words and formatting make this vary considerably. This is why we use a character-chunk size of 800800—to make sure we have sufficient head-room to always end up with a $100$ token sequence.) We then shift $10$ characters forward in the book text and repeat this process. We do this for the entire length of the book, which results in approximately one sequence every $10$ characters, i.e., $\texttt{len(book\_characters)}/10$ total sequences $\bm{z}$}, for which we compute $p_{\bm{z}}$. For example, The Great Gatsby has 270,870270{,}870 characters, which results in roughly 27,00027{,}000 sequences, for which we compute $p_{\bm{z}}$ (Equation A.2). (See Section 2 and Appendix A.2.)This means that the $100$-token sequences overlap significantly. This is deliberate. We typically do not know how different open-weight models were trained; a priori, it is not clear exactly where we should begin sequences (or where they should end). With this approach, for extractable content, we expect to surface high-probability regions of memorized content, which we can then explore in more detail and more precisely in follow-up experiments. We picked a sliding window of $10$ characters by running initial experiments with $1$-, $10$-, $50$-, and $100$-character windows. $10$ characters exhibited an insignificant loss in extraction signal (at $10\times$ cheaper cost than $1$ character). We intend to investigate reducing this cost further in the future. Initially, we did not expect this approach to result in high per-book extraction rates. This is because, even though this approach should surface memorization “hot-spots” by getting good coverage of $100$-token sequences across the book, we are not being particularly clever about the start location for each sequence. Since we indiscriminately slide the window $10$ characters each time, we expect that many prompts will begin in “odd” locations—locations in the middle of tokens, which lead to “broken” or otherwise “strange” tokenizations. We expected to yield near-zero probability of extraction.However, with the final presentation of our results—in which we use a very high minimum probability $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$ in this work for deeming extraction successful (Sections 2.15; Appendices CE)—this did not really have an impact on extraction rates.So, in the subsections that follow, we show per-book extraction rates across the different models we test with this procedure. These are useful for very high-level comparisons of how many sequences we extracted with probabilistic extraction, compared to greedy-decoded discoverable extraction as a baseline (which, similar to Hayes et al. [60], we observe consistently under-counts extraction).
@@
Similar to Hayes et al. [60], after some initial tests with different decoding schemes, we use temperature T=1T{=}1 for top-$k$ decoding with the k=40k{=}40. We believe that setting T=1T{=}1 makes the most sense for studying memorization, as this reflects the LLM’s base probability distribution. Since extraction is only successful when the underlying training sequence’s logits reflect very high probabilities (Section 2), we also think it makes sense to clip the full-vocabulary logit distribution to reflect some number of the top-highest probability tokens per step. This is the same intuition behind using greedy decoding for discoverable extraction—but with a twist. Greedy decoding deterministically picks the single-most-likely token per step. (This is equivalent to top-$k$ decoding with T=1T{=}1 and k=1k{=}1.) As the results in Hayes et al. [60] show, this can be a suboptimal decoding strategy for measuring extraction, for several reasons:
@@
  • More generally, this allows for a bit more “wiggle room.” By being able to pick beyond just the top-$1$ token per step, the decoding process take locally suboptimal steps, which can still result in high-probability sequences that (even if slightly lower probability than the greedy sequence, with respect to the normalized top-$k$ distribution) reflect extraction. While we do not investigate this in detail in this work, we observe that (when this happens) such sequences are very similar to the greedy-decoded suffix; the greedy-decoded suffix is a near-miss for the verbatim suffix (it is near-verbatim). It may be reasonable under more relaxed (near-verbatim) metrics for extraction to have counted this greedy suffix as extracted, but that is not the approach we (or Hayes et al. [60]) take, and so these suffixes would be missed. We intend to investigate this further in future work.
@@
The above reasons also reflect the general motivations for using probabilistic extraction (rather than traditional discoverable extraction) as our metric (Sections 23; Appendix A). Similar motivations for useful extraction measurements are also noted in older extraction papers (e.g., Carlini et al. [16], Carlini et al. [18]); however, this work typically opts to use greedy decoding based on the perception that it is cheaper to do so at scale. This work does not claim that greedy decoding is the best or only way to measure extraction, and in fact notes [16]—and in select cases shows [18]—that other (more expensive) decoding schemes are able to extract more training data. (Again, this seems likely because these schemes, like beam search [18], are able to produce higher-likelihood sequences under the model.)Further, as we show through optimizing the work of Hayes et al. [60], it turns out that it is actually much cheaper to compute probabilistic extraction (with any top-$k$ decoding configuration, including greedy/ k=1k{=}1), as we compute extraction probabilities as a post-processing operation directly off of the logits obtained from inference, rather than (as has been done in prior work [16, 18]) actually generating (greedy) completions. We can compute verbatim extraction probabilities with a single forward pass through the model, which is an enormous cost savings compared to generating $50$ tokens per prompt. We discuss these details in Appendix A.3. Note that because our implementation of probabilistic extraction makes computing $\log$ probabilities a post-processing operation on the output model logits on CPU (Appendix A), we also collect results for greedy-decoded discoverable extraction [74, 16, 93] by setting T=k=1T{=}k{=}1, incurring negligible cost to do so. For brevity, we do not include detailed results for this setting. To give a brief intuition for these results, we do include greedy extraction rates (with respect to the set of sliding-window sequences we construct) for each book’s results in Appendix D.3; these rates consistently under-count probabilistic extraction with top-$k$ decoding (T=1T{=}1, k=40k{=}40).
@@ D.3 Book-specific results
  • Bar plot comparing extraction and generation rates. For each model, we plot three different rates: the greedy-decoded discoverable extraction rate; the probabilistic extraction rate for top-$k$ decoding (T=1T{=}1, k=40k{=}40), counting a sequence as extracted if $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$; and the maximum possible rate of generating suffixes given their corresponding prefixes. Each rate is computed over all (overlapping) sequences from the sliding-window procedure (Appendix D.2). Because we use a consistent sequence set, this comparison is valid even with overlaps. The discoverable extraction rate consistently underestimates the probabilistic extraction rate, even with this conservative $\tau_{\text{min}}$. The maximum rate of suffix generation is not an extraction rate: here we do not apply $\tau_{\text{min}}$ (effectively setting it to $0\%$), so it includes all sequences with nonzero $p_{\bm{z}}$. This captures not only valid extractions but also low-probability instances of spontaneous suffix generation given a prefix. We include this number for two reasons:
@@
      It is never $100\%$, showing that not every suffix can be generated given its prefix, even in the limit of repeated prompts. This is important to emphasize for legal audiences, where there is a common misconception that, with enough queries, a probabilistic model can generate anything [33]. Our results show that this is not the case.
@@
If no sequences are extracted for a given model–book pair, we omit histograms (they would be empty). In one case—The Beautiful Struggle [22]—we extract nothing for any model, so both histograms and heatmaps are omitted. For histograms and heatmaps, results are presented in three groups corresponding to the categories of models described in Appendix D.1: models known with certainty to have been trained on Books3 (blue), models that were likely trained on books also present in Books3 (red), and one model (Phi 4) that, based on documentation, was not trained on whole copyrighted books (orange). See Section 5 and Appendix E for further discussion about how we can be confident that our results for all three categories reflect valid extraction.
@@
Visualizing memorization “hot-spots” also allows us to identify runs of memorized text from which we can attempt to extract more than $50$ tokens. That is, in regions with very high extraction probabilities, these probabilities may be high enough to yield nontrivial extraction probabilities for longer runs. We take the full list of $p_{\bm{z}}$ for all sequences in a given book–model experiment, along with their start and end positions, sorted by start position. We filter to only those with especially high probability (e.g., $p_{\bm{z}}\geq 50\%$), then merge overlapping sequences by taking the minimum start and maximum end. From this we construct longer candidate sequences to attempt extraction. To illustrate why this works, consider a single $50$-token suffix with $p_{\bm{z}}=50\%$. Its per-token geometric mean probability is $0.5^{1/50}\approx 98.62\%$. For this individual token probability, the longest sequence length $x$ that maintains $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$ is
@@
(However, for a suffix of this length, it would perhaps be reasonable to consider a smaller $\tau_{\text{min}}$ for determining extraction success.) If we see multiple sequences close by that exhibit $p_{\bm{z}}=50\%$, we could perhaps find a much longer sequence that fits the conditions above (or something similar). We applied this methodology to Harry Potter and the Sorcerer’s Stone [106] and 1984 [100] with Llama 3.1 70B, given the extent of memorization we observe in those cases. Figure 13 shows one successful case: a $300$-token extraction consisting of a $25$-token prefix and a $275$-token suffix with $p_{\bm{z}}=36.1\%$.
Figure 13: An example of a long sequence from Harry Potter and the Sorcerer’s Stone [106] that we extract from Llama 3.1 70B, with top-$k$ (T=1T{=}1 k=40k{=}40) decoding. The prompt is a $25$-token prefix; the target suffix is $275$ tokens. This suffix has $p_{\bm{z}}=36.1\%$, which corresponds to a geometric-mean individual token conditional probability of over $99.6\%$, i.e., $0.9963^{275}\approx 36.1\%$). In expectation, with this prompt, this suffix is generated verbatim more than $\nicefrac{{1}}{{3}}$ of the time.
D.3.1 The Handmaid’s Tale, Atwood
@@
Figure 14: The Handmaid’s Tale, Atwood [5]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.2 Dante and the Origins of Italian Culture, Barolini
@@
Figure 16: Dante and the Origins of Italian Culture, Barolini [7]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.3 The Da Vinci Code, Brown
@@
Figure 18: The Da Vinci Code, Brown [11]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.4 The Myth of Sisyphus, Camus

### v4

@@ D.2 Sliding-window probabilistic extraction procedure
We use $100$-token sequences $\bm{z}$}, with a prefix (prompt) length of $50$ tokens and a (target) suffix length of $50$ tokens. We pick $100$ tokens and a 5050/$50$ split because this is accepted in the literature as a reasonable minimum for being confident that extracted suffixes are reflective of memorization [carlini2023quantifying, carlini2021extracting, lee2022dedup, hayes2025measuringmemorizationlanguagemodels]. $50$ tokens for the prefix is less for validity reasons. Of course, this prefix length uses an amount of context that is equivalent to the output, and so is compelling in this respect. However, when we are confident that our measurements are capturing valid instances of extraction, using longer prefixes can help discover more underlying memorization. (Indeed, this is part of where discoverable extraction gets its name; carlini2023quantifying demonstrate the “discoverability phenomenon” with respect to showing how extraction success can increase with longer-context prompts.) We discuss this further (and include experiments with varied prefix lengths) in Section 5 and Appendix E. Other work uses shorter sequences (e.g., $64$ tokens and a 3232/$32$ token split [biderman2023pythia]).
@@
For a given book, we start at the beginning of the text file in Books3. We sample a chunk of text that is sufficiently long to contain $100$ tokens of corresponding tokenized text: to do this, we take a chunk of $800$ characters, tokenize, take the first $100$ tokens of the resulting tokenized sequence as the sequence, and discard the rest. (Each $100$ token sequence is typically, but not always, reflective of 300300-$400$ characters of text. Rare words and formatting make this vary considerably. This is why we use a character-chunk size of 800800—to make sure we have sufficient head-room to always end up with a $100$ token sequence.) We then shift $10$ characters forward in the book text and repeat this process. We do this for the entire length of the book, which results in approximately one sequence every $10$ characters, i.e., $\texttt{len(book\_characters)}/10$ total sequences $\bm{z}$}, for which we compute $p_{\bm{z}}$. For example, The Great Gatsby has 270,870270{,}870 characters, which results in roughly 27,00027{,}000 sequences, for which we compute $p_{\bm{z}}$ (Equation A.2). (See Section 2 and Appendix A.2.)
This means that the $100$-token sequences overlap significantly. This is deliberate. We typically do not know how different open-weight models were trained; a priori, it is not clear exactly where we should begin sequences (or where they should end). With this approach, for extractable content, we expect to surface high-probability regions of memorized content, which we can then explore in more detail and more precisely in follow-up experiments. We picked a sliding window of $10$ characters by running initial experiments with $1$-, $10$-, $50$-, and $100$-character windows. $10$ characters exhibited an insignificant loss in extraction signal (at $10\times$ cheaper cost than $1$ character). We intend to investigate reducing this cost further in the future.
Initially, we did not expect this approach to result in high per-book extraction rates. This is because, even though this approach should surface memorization “hot-spots” by getting good coverage of $100$-token sequences across the book, we are not being particularly clever about the start location for each sequence. Since we indiscriminately slide the window $10$ characters each time, we expect that many prompts will begin in “odd” locations—locations in the middle of tokens, which lead to “broken” or otherwise “strange” tokenizations. We expected to yield near-zero probability of extraction.
However, with the final presentation of our results—in which we use a very high minimum probability $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$ in this work for deeming extraction successful (Sections 2.1 & 5; Appendices C & E)—this did not really have an impact on extraction rates.
So, in the subsections that follow, we show per-book extraction rates across the different models we test with this procedure. These are useful for very high-level comparisons of how many sequences we extracted with probabilistic extraction, compared to greedy-decoded discoverable extraction as a baseline (which, similar to hayes2025measuringmemorizationlanguagemodels, we observe consistently under-counts extraction).
@@
Similar to hayes2025measuringmemorizationlanguagemodels, after some initial tests with different decoding schemes, we use temperature T=1T{=}1 for top-$k$ decoding with the k=40k{=}40. We believe that setting T=1T{=}1 makes the most sense for studying memorization, as this reflects the LLM’s base probability distribution. Since extraction is only successful when the underlying training sequence’s logits reflect very high probabilities (Section 2), we also think it makes sense to clip the full-vocabulary logit distribution to reflect some number of the top-highest probability tokens per step. This is the same intuition behind using greedy decoding for discoverable extraction—but with a twist. Greedy decoding deterministically picks the single-most-likely token per step. (This is equivalent to top-$k$ decoding with T=1T{=}1 and k=1k{=}1.) As the results in hayes2025measuringmemorizationlanguagemodels show, this can be a suboptimal decoding strategy for measuring extraction, for several reasons:
@@
  • More generally, this allows for a bit more “wiggle room.” By being able to pick beyond just the top-$1$ token per step, the decoding process take locally suboptimal steps, which can still result in high-probability sequences that (even if slightly lower probability than the greedy sequence, with respect to the normalized top-$k$ distribution) reflect extraction. While we do not investigate this in detail in this work, we observe that (when this happens) such sequences are very similar to the greedy-decoded suffix; the greedy-decoded suffix is a near-miss for the verbatim suffix (it is near-verbatim). It may be reasonable under more relaxed (near-verbatim) metrics for extraction to have counted this greedy suffix as extracted, but that is not the approach we (or hayes2025measuringmemorizationlanguagemodels) take, and so these suffixes would be missed. We intend to investigate this further in future work.
@@
The above reasons also reflect the general motivations for using probabilistic extraction (rather than traditional discoverable extraction) as our metric (Sections 2 & 3; Appendix A). Similar motivations for useful extraction measurements are also noted in older extraction papers (e.g., carlini2021extracting, carlini2023quantifying); however, this work typically opts to use greedy decoding based on the perception that it is cheaper to do so at scale. This work does not claim that greedy decoding is the best or only way to measure extraction, and in fact notes [carlini2021extracting]—and in select cases shows [carlini2023quantifying]—that other (more expensive) decoding schemes are able to extract more training data. (Again, this seems likely because these schemes, like beam search [carlini2023quantifying], are able to produce higher-likelihood sequences under the model.)
Further, as we show through optimizing the work of hayes2025measuringmemorizationlanguagemodels, it turns out that it is actually much cheaper to compute probabilistic extraction (with any top-$k$ decoding configuration, including greedy/ k=1k{=}1), as we compute extraction probabilities as a post-processing operation directly off of the logits obtained from inference, rather than (as has been done in prior work [carlini2021extracting, carlini2023quantifying]) actually generating (greedy) completions. We can compute verbatim extraction probabilities with a single forward pass through the model, which is an enormous cost savings compared to generating $50$ tokens per prompt. We discuss these details in Appendix A.3.
Note that because our implementation of probabilistic extraction makes computing $\log$ probabilities a post-processing operation on the output model logits on CPU (Appendix A), we also collect results for greedy-decoded discoverable extraction [lee2022dedup, carlini2021extracting, nasr2023scalable] by setting T=k=1T{=}k{=}1, incurring negligible cost to do so. For brevity, we do not include detailed results for this setting. To give a brief intuition for these results, we do include greedy extraction rates (with respect to the set of sliding-window sequences we construct) for each book’s results in Appendix D.3; these rates consistently under-count probabilistic extraction with top-$k$ decoding (T=1T{=}1, k=40k{=}40).
@@ D.3 Book-specific results
  • Bar plot comparing extraction and generation rates. For each model, we plot three different rates: the greedy-decoded discoverable extraction rate; the probabilistic extraction rate for top-$k$ decoding (T=1T{=}1, k=40k{=}40), counting a sequence as extracted if $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$; and the maximum possible rate of generating suffixes given their corresponding prefixes. Each rate is computed over all (overlapping) sequences from the sliding-window procedure (Appendix D.2). Because we use a consistent sequence set, this comparison is valid even with overlaps.
    The discoverable extraction rate consistently underestimates the probabilistic extraction rate, even with this conservative $\tau_{\text{min}}$. The maximum rate of suffix generation is not an extraction rate: here we do not apply $\tau_{\text{min}}$ (effectively setting it to $0\%$), so it includes all sequences with nonzero $p_{\bm{z}}$. This captures not only valid extractions but also low-probability instances of spontaneous suffix generation given a prefix. We include this number for two reasons:
@@
      It is never $100\%$, showing that not every suffix can be generated given its prefix, even in the limit of repeated prompts. This is important to emphasize for legal audiences, where there is a common misconception that, with enough queries, a probabilistic model can generate anything [toothpick]. Our results show that this is not the case.
@@
If no sequences are extracted for a given model–book pair, we omit histograms (they would be empty). In one case—The Beautiful Struggle [struggle]—we extract nothing for any model, so both histograms and heatmaps are omitted. For histograms and heatmaps, results are presented in three groups corresponding to the categories of models described in Appendix D.1: models known with certainty to have been trained on Books3 (blue), models that were likely trained on books also present in Books3 (red), and one model (Phi 4) that, based on documentation, was not trained on whole copyrighted books (orange). See Section 5 and Appendix E for further discussion about how we can be confident that our results for all three categories reflect valid extraction.
@@
Visualizing memorization “hot-spots” also allows us to identify runs of memorized text from which we can attempt to extract more than $50$ tokens. That is, in regions with very high extraction probabilities, these probabilities may be high enough to yield nontrivial extraction probabilities for longer runs.
We take the full list of $p_{\bm{z}}$ for all sequences in a given book–model experiment, along with their start and end positions, sorted by start position. We filter to only those with especially high probability (e.g., $p_{\bm{z}}\geq 50\%$), then merge overlapping sequences by taking the minimum start and maximum end. From this we construct longer candidate sequences to attempt extraction.
To illustrate why this works, consider a single $50$-token suffix with $p_{\bm{z}}=50\%$. Its per-token geometric mean probability is $0.5^{1/50}\approx 98.62\%$. For this individual token probability, the longest sequence length $x$ that maintains $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$ is
@@
(However, for a suffix of this length, it would perhaps be reasonable to consider a smaller $\tau_{\text{min}}$ for determining extraction success.) If we see multiple sequences close by that exhibit $p_{\bm{z}}=50\%$, we could perhaps find a much longer sequence that fits the conditions above (or something similar).
We applied this methodology to Harry Potter and the Sorcerer’s Stone [hp1] and 1984 [1984] with Llama 3.1 70B, given the extent of memorization we observe in those cases. Figure 13 shows one successful case: a $300$-token extraction consisting of a $25$-token prefix and a $275$-token suffix with $p_{\bm{z}}=36.1\%$.
Figure 13: An example of a long sequence from Harry Potter and the Sorcerer’s Stone [hp1] that we extract from Llama 3.1 70B, with top-$k$ (T=1T{=}1 k=40k{=}40) decoding. The prompt is a $25$-token prefix; the target suffix is $275$ tokens. This suffix has $p_{\bm{z}}=36.1\%$, which corresponds to a geometric-mean individual token conditional probability of over $99.6\%$, i.e., $0.9963^{275}\approx 36.1\%$). In expectation, with this prompt, this suffix is generated verbatim more than $\nicefrac{{1}}{{3}}$ of the time.
D.3.1 The Handmaid’s Tale, handmaid
@@
Figure 14: The Handmaid’s Tale, handmaid: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.2 Dante and the Origins of Italian Culture, dante
@@
Figure 16: Dante and the Origins of Italian Culture, dante: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.3 The Da Vinci Code, davinci
@@
Figure 18: The Da Vinci Code, davinci: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.4 The Myth of Sisyphus, sisyphus